# SAM-CP: Marrying SAM with Composable Prompts for Versatile Segmentation

### v1

@@ 4 Experiments
PQ SQ RQ AP mIoU PQ SQ RQ AP mIoU PQ AP mIoU
MaskCLIP [15] VIT-L 15.1 70.5 19.2 6.0 23.7
FreeSeg [39] VIT-B 16.3 71.8 21.6 6.5 24.6 21.7 72.0 21.6 6.6 21.7
ODISE [52] VIT-H 23.3 74.4 27.9 13.0 29.2 25.0 79.4 30.4 23.9
OPSNet [10] VIT-L 19.0 52.4 23.0 41.5
MaskQCLIP [53] VIT-L 23.3 30.4
X-Decoder [63] Focal-L 21.8 13.1 29.6 38.1 24.9 52.0
FCCLIP [56] CN-L 26.8 71.5 32.3 16.8 34.1 27.0 78.0 32.9 44.0 26.8 56.2
SAM-CP CN-L 27.2 77.7 32.9 17.0 31.8 28.6 78.4 34.5 21.9 34.3 41.0 29.3 47.9
@@
DETR[3] R50 reg.+seg. 50+25e 31.1
MaskFormer [12] R50 seg. 300e 46.5 33.9 57.8 128e 34.7
Mask2Former [11] R50 seg. 50e 51.5 41.7 61.7 128e 39.7 26.4 47.7
Mask2Former [11] Swin-L seg. 100e 57.8 48.6 67.4 128e 48.1 34.2 56.1
Mask DINO [27] R50 reg.+seg. 50e 53.0 48.8 44.3 60.6
Mask DINO [27] Swin-L reg.+seg. 50e 58.3 56.2 50.6 67.3 128e 56.6
X-Decoder [63] Focal-T seg. 50e 52.6 41.3 62.4 128e 41.6 27.7 51.0
X-Decoder [63] Focal-L seg. 50e 56.9 46.7 67.5 128e 49.6 35.8 58.1
@@ 4.1 Settings
Datasets and evaluation metrics. We train SAM-CP on the COCO-Panoptic [32] and ADE20K [61] datasets, and evaluate the models on either closed-domain or open-vocabulary segmentation (with cross-dataset validation and Cityscapes used as test data). COCO-Panoptic (the 2017 version) has 118K training and 5K validation images with $80$ ‘thing’ and $53$ ‘stuff’ categories. We report the instance segmentation results with the standard AP metric on the $80$ ‘thing’ categories. For semantic segmentation, we report the mIoU for all ($80+53$) categories. For panoptic segmentation, the PQ metric is computed for all categories and the ‘thing’ and ‘stuff’ subsets individually. ADE20K contains 20,210 images. We use the $150$ most common object categories, including $100$ ‘thing’ and $50$ ‘stuff’ categories. Cityscapes is a street view dataset with 8 ‘thing’ categories and 11 ‘stuff’ categories. We inherit the same metric from COCO to ADE20K and Cityscapes datasets.
Implementation details. The proposal masks are generated by SAM with $48$ grid points along each axis of the input image. To verify our idea better, we use SAM with VIT-H to generate better patches. For open-vocabulary segmentation, we use a frozen CLIP image encoder (for a fair comparison, we use the same architecture (ConvNext-L) as the previous best method, FCCLIP [56]) as the backbone and equip it with FPN [30]. For closed-domain segmentation, we establish SAM-CP on ResNet50 (R50) [20] and Swin-L [36]. We use the implementation of the MMDetection [7] (v3.0) library. We use $8$ Tesla-V100 GPUs ($4$/$2$ images per GPU) for open-vocabulary/closed-domain experiments. The data augmentation strategy follows the DETR series. An AdamW optimizer [37] with a basic learning rate of $0.0002$. See Appendix B for further details.
@@ 4.2 Quantitative Results
Closed-domain segmentation. Results are summarized in Table 2. In COCO-Panoptic, with the ResNet-50 backbone, SAM-CP achieves a $48.6\%$ PQ, $41.7\%$ AP, and $55.6\%$ mIoU; with a stronger backbone, Swin-L, SAM-CP reports higher segmentation accuracy, with a $52.7\%$ PQ, $45.2\%$ AP, and $61.7\%$ mIoU. In ADE20K, SAM-CP achieves a $38.5\%$ PQ, a $25.1\%$ AP, and a $42.4\%$ mIoU on ResNet-50, and a $44.4\%$ PQ, a $30.3\%$ AP, and a $49.4\%$ mIoU on Swin-L. An interesting comparison comes from MaskFormer [12] where SAM-CP reports higher PQ and AP but a lower mIoU, which implies its advantageous performance in instance-level recognition. These numbers demonstrate the effectiveness of our methodology, i.e., establishing composable prompts beyond vision foundation models. We will delve into the limitation of SAM-CP in Section 4.5 and explain why this novel mechanism falls short in closed-domain segmentation.
@@ 4.4 Ablative Studies
Adding extra proposals. We add the proposals extracted by the pre-trained Mask DINO (MD) model into the pool of candidate patches. Details are provided in Appendix B.1. Table 10 shows improved segmentation results, inspiring us that SAM does not generate ideal patches, especially for closed-set segmentation, and we look forward to stronger vision foundation models in the future.
Figure 5: A showcase of our results on COCO for closed domain. In each group, from left to right: input, SAM patches, panoptic segmentation by SAM-CP (ours). This figure is best viewed in color.
$\tau$ PQ APdet AP mIoU
0.9 43.2 42.8 38.2 49.9
0.8 45.0 45.4 40.7 52.6
0.7 43.3 43.3 39.9 50.3
0.5 42.0 34.3 37.2 48.9
0.8 (w/o $\mathrm{IoP}_{\mathrm{box}}$) 42.1 31.7 35.4 48.4
0.8 (w/o $\mathrm{IoP}_{\mathrm{mask}}$) 39.7 42.3 32.5 44.1
0.8 (w/low quality match) 47.0 45.8 41.4 54.2
@@
4.5 Advantages and Limitations
SAM-CP inherits a clear advantage from SAM in the generalized ability across different visual domains. SAM-CP achieves this by decomposing low-level pixel grouping (offered by SAM) from high-level semantic recognition (offered by the composable prompts). In the meantime, the imperfection of SAM limits the segmentation accuracy of SAM-CP, especially in the closed-set vision benchmarks. We showcase this point in Table 10 where the patches generated by SAM are compared to those generated by Mask DINO [27], a state-of-the-art segmentation model. We find that SAM suffers a higher missing rate (e.g., one cannot find a proposal with an IoU larger than $0.5$ for $16.7\%$ instances; the rate is not significantly smaller even if we refer to the ground-truth masks to merge some proposals) than Mask DINO (where the rate is only $10.1\%$). In other words, SAM failed to find some objects or incorrectly merged two or more objects into one patch, and SAM-CP cannot save such loss. This contributes to the deficit in the segmentation accuracy (e.g., AP or mIoU) compared to Mask DINO. We display some typical examples in Appendix A.4. Additionally, the inference speed of SAM-CP is bound by that of SAM; once a more efficient vision foundation model is available, our framework can be seamlessly transplanted and achieve faster inference.
@@ A.1 Affinity Computation
Dynamic cross-attention. Rather than performing global computation, we use dynamic cross-attention (DCA) to guide the cross-attention operator in focusing on the local features. This is motivated by deformable attention [62]. The cross-attention operator follows the same structure of multi-head attention [3], and the matrix $\mathbf{A}$ serves as the dynamic attention mask. We binarize $\mathbf{A}$ by setting the value of each entry to $1$ if the value is smaller than the threshold (e.g., $0.5$) which means that the area is masked and ignored in computing the cross-attention. As described in Section 3.2.2, the query vectors (as Q) and the patch features (as both K and V) are fed into a multi-head cross-attention with the dynamic attention mask.
@@ A.2 Supervision
Ground-truth class labels. Let $c_{m}^{\star}$ be the ground-truth class label for the $m$-th query. For a Type-I prompt, the $m$-th query corresponds to the $c_{m}^{\star}$-th category label. Hence, the objective of the $m$-th query is set to be $c_{m}^{\star}$ if the image has the $c_{m}^{\star}$-th category label, and ‘negative’ otherwise. For a Type-II prompt, $c_{m}^{\star}$ comes from the label assignment of the Hungarian matching algorithm. Hence, the objective of the $m$-th query is set to be $c_{m}^{\star}$ if it is a positive (matched) query, and ‘negative’ otherwise.
@@
Figure 6: The illustration of how to get the ground-truth affinity matrix $\mathbf{B}$.
A.3 The Denoising Strategy
We follow DINO [59] to enhance the stability of optimization. We increase some queries generated by adding some noise for the ground-truth instances and supervise them using the corresponding ground truth in the training stage. The denoising strategy can make the training more stable because one ground truth may be matched to one query in this epoch but be matched to another in the next epoch, which is unfavourable for optimization.
For each instance, we shake the bounding box of the object as the position embedding and provide incorrect category features as the query vectors in a certain proportion, which is the same as DINO. Then, the denoising queries are concatenated to the original queries for prediction. Note that the original queries cannot see these queries in self-attention because the denoising queries will be removed in the inference stage. Then, in loss calculation, the denoising queries do not participate in label assignment and are directly supervised by the corresponding class and affinity matrix labels. We refer the readers to the original DINO paper for further details.
[t] Refer to caption
@@ A.3 The Denoising Strategy
A.4 Examples for Advantages and Limitations
@@
Data preparation. We use the same data augmentation strategy in DETR-based detector [3, 62, 59]. We randomly flip the image, resize the image ranging from $1333\times 480$ to $1333\times 800$, randomly crop the image, and randomly resize the image again. In the inference stage, we resize the image to $1333\times 800$ without any other augmentation. For SAM proposal generation, we use the everything mode in SAM with $48$ points per side, a $0.8$ IoU threshold, a $0.9$ stability score threshold, and a $0.7$ NMS threshold.
@@
Add patches from Mask DINO. We add the patches from Mask DINO [27] to show that the upper bound of SAM patches limits our performance. If the quality of patches generated by the foundation model gets better in the future, the segmentation accuracy of SAM-CP will further increase. We saved the segmentation results of Mask DINO and deleted the label result and confidence score. Then, the segmentation results are viewed as the extra proposal and added to the SAM patches.
@@
Data preparation. We use the same data augmentation strategy in Mask2former [11] in ADE20K dataset. The image is randomly flipped and resized in the range of $[320,1280]$. Then, the image is randomly cropped by $2560\times 640$. In the inference stage, the image is resized to $2560\times 640$. The hyper-parameters for SAM patch generation are the same as those for the COCO dataset.
@@
Setting for the model. The training hyper-parameters are the same as those in close-domain experiments. Only the trainable backbone is replaced by a frozen CLIP image encoder. In the inference stage, the parameter $\kappa$ described in Section 3.3 is set as 0.4. In addition, we follow FCCLIP [56] to use the prompt engineering when inference. Each label can derive multiple entries, and we chose the highest score as the classification score for this label.
@@
We provide our codes in https://github.com/paperrepo123/SAM-CP. After the acceptance of our paper, we will make our codes publically available.
@@
Appendix C More Visualization Results
C.1 More t-SNE Visualization
Figure 9 shows more t-SNE visualizations of our approach. The features of patches belong to different semantic/instance groups to different clusters in SAM-CP. Not that, in lines 6–10, the different instances with the same label (’train’, ’elephant’, ’person’, ’horse’, and ’umbrella’) cluster together to the same semantic area but separate to different instances in the t-SNE feature space.
C.2 More Segmentation Examples

### v2

@@ 4 Experiments
PQ SQ RQ AP mIoU PQ SQ RQ AP mIoU AP PQ mIoU
MaskCLIP Ding et al. (2023) VIT-L 15.1 70.5 19.2 6.0 23.7
FreeSeg Qin et al. (2023) VIT-B 16.3 71.8 21.6 6.5 24.6 21.7 72.0 21.6 6.6 21.7
OPSNet Chen et al. (2023c) VIT-L 19.0 52.4 23.0 41.5
OpenSeeD Zhang et al. (2023b) Focal-L 19.7 15.0 29.0 41.4
HIPIE Wang et al. (2023d) VIT-H 22.9 19.0 29.0
X-Decoder Zou et al. (2023) Focal-L 21.8 13.1 29.6 24.9 38.1 52.0
MaskQCLIP Xu et al. (2023b) VIT-L 23.3 30.4
ODISE Xu et al. (2023a) VIT-H 23.3 74.4 27.9 13.0 29.2 25.0 79.4 30.4 23.9
CLIPSelf Wu et al. (2024) VIT-L 23.7 13.6 30.1
FrozenSeg Chen et al. (2024b) CN-L 25.9 16.4 34.4 28.4 45.8 56.8
FCCLIP Yu et al. (2023a) CN-L 26.8 71.5 32.3 16.8 34.1 27.0 78.0 32.9 26.8 44.0 56.2
SAM-CP CN-L 27.2 77.7 32.9 17.0 31.8 28.6 78.4 34.5 21.9 34.3 29.3 41.0 47.9
@@
DETR Carion et al. (2020) R50 reg.+seg. 50+25e 31.1
MaskFormer  Cheng et al. (2021) R50 seg. 300e 46.5 33.9 57.8 128e 34.7
Mask2Former  Cheng et al. (2022) R50 seg. 50e 51.5 41.7 61.7 128e 39.7 26.4 47.7
Mask2Former  Cheng et al. (2022) Swin-L seg. 100e 57.8 48.6 67.4 128e 48.1 34.2 56.1
Mask DINO  Li et al. (2023b) R50 reg.+seg. 50e 53.0 48.8 44.3 60.6
Mask DINO  Li et al. (2023b) Swin-L reg.+seg. 50e 58.3 56.2 50.6 67.3 128e 56.6
X-Decoder  Zou et al. (2023) Focal-T seg. 50e 52.6 41.3 62.4 128e 41.6 27.7 51.0
X-Decoder  Zou et al. (2023) Focal-L seg. 50e 56.9 46.7 67.5 128e 49.6 35.8 58.1
@@
4.1 Inference
The inference procedure varies slightly between closed-set and open-vocabulary segmentation. In a closed domain, the logit $\mathbf{S}^{\mathrm{cls}}$ in Equation 1 is used for classification. The normalized and quantized affinity matrix $\mathbf{A}$ is used for patch merging. For semantic segmentation, we refer to the rows corresponding to $C$ classes and merge all patches surpassing a pre-defined threshold. For instance segmentation, we look up the non-empty instance rows of $\mathbf{A}$, each of which corresponds to an instance. Panoptic segmentation is achieved by combining semantic and instance segmentation results. In an open domain, we complement the logit matrix $\mathbf{S}^{\mathrm{cls}}$ (sized $M\times C$) with a CLIP-based logit matrix, $\mathbf{S}^{\mathrm{CLIP}}$. The remaining part is the same as in a closed domain. To calculate $\mathrm{S}^{\mathrm{CLIP}}$, we follow FC-CLIP Yu et al. (2023a) to extract CLIP features using mask pooling on the predicted masks. Then, the feature $\mathbf{S}^{\mathrm{CLIP}}$ is obtained by calculating the similarity between the CLIP feature and $\hat{\mathbf{e}}_{c}$. The final class score is computed as $\mathbf{S}^{\mathrm{ov}}=(\sigma(\mathbf{S}^{\mathrm{cls}}))^{(1-\kappa)}+(\mathrm{softmax}(\mathbf{S}^{\mathrm{CLIP}}))^{\kappa}$, where $\kappa=0.4$ is a coefficient balancing the closed-domain and open-vocabulary class scores.
Datasets and evaluation metrics. We train SAM-CP on the COCO-Panoptic Lin & Maire (2014) and ADE20K Zhou et al. (2017) datasets, and evaluate the models on either closed-domain or open-vocabulary segmentation (with cross-dataset validation and Cityscapes used as test data). COCO-Panoptic (the 2017 version) has 118K training and 5K validation images with $80$ ‘thing’ and $53$ ‘stuff’ categories. We report the instance segmentation results with the standard AP metric on the $80$ ‘thing’ categories. For semantic segmentation, we report the mIoU for all ($80+53$) categories. For panoptic segmentation, the PQ metric is computed for all categories and the ‘thing’ and ‘stuff’ subsets individually. ADE20K contains 20,210 images. We use the $150$ most common object categories, including $100$ ‘thing’ and $50$ ‘stuff’ categories. Cityscapes is a street-view dataset with 8 ‘thing’ and 11 ‘stuff’ categories. We inherit the same metric from COCO to ADE20K and Cityscapes datasets.
@@ 4.2 Settings
Implementation details. The proposal masks are generated by SAM with $48$ grid points along each axis of the input image. To verify our idea better, we use SAM with VIT-H to generate better patches. For open-vocabulary segmentation, we use a frozen CLIP image encoder (for a fair comparison, we use the same architecture (ConvNext-L) as the previous best method, FCCLIP Yu et al. (2023a)) as the backbone and equip it with FPN Lin et al. (2017a). For closed-domain segmentation, we establish SAM-CP on ResNet50 (R50) He et al. (2016) and Swin-L Liu et al. (2021). We use the implementation of the MMDetection Chen et al. (2019) (v3.0) library. We use $8$ Tesla-V100 GPUs ($4$/$2$ images per GPU) for open-vocabulary/closed-domain experiments. The data augmentation strategy follows the DETR series. An AdamW optimizer Loshchilov & Hutter (2019) with a basic learning rate of $0.0002$. See Appendix B for further details.
@@ 4.3 Quantitative Results
Closed-domain segmentation. Results are summarized in Table 2. In COCO-Panoptic, with the ResNet-50 backbone, SAM-CP achieves a $48.6\%$ PQ, $41.7\%$ AP, and $55.6\%$ mIoU; with a stronger backbone, Swin-L, SAM-CP reports higher segmentation accuracy, with a $52.7\%$ PQ, $45.2\%$ AP, and $61.7\%$ mIoU. In ADE20K, SAM-CP achieves a $38.5\%$ PQ, a $25.1\%$ AP, and a $42.4\%$ mIoU on ResNet-50, and a $44.4\%$ PQ, a $30.3\%$ AP, and a $49.4\%$ mIoU on Swin-L. An interesting comparison comes from MaskFormer Cheng et al. (2021) where SAM-CP reports higher PQ and AP but a lower mIoU, which implies its advantageous performance in instance-level recognition. These numbers demonstrate the effectiveness of our methodology, i.e., establishing composable prompts beyond vision foundation models. We will delve into the limitation of SAM-CP in Section 4.6 and explain why this novel mechanism falls short in closed-domain segmentation.
@@ 4.5 Ablative Studies
Method mAP $AP_{50}$ $AP_{75}$
w.o. general 13.7 30.5 10.6
w. general 13.6 30.1 10.5
Table 10: Accuracy (%) of part segmentation on Pascal Part dataset. The first row shows results of SAM-CP trained only on part datasets, while the second shows those of SAM-CP trained on both part and general datasets.
Method SAM - based SAM / non-SAM time (s/img) total time (s/img) PQ AP mIoU
FCCLIP Yu et al. (2023a) no 0 / 0.24 0.24 25.2 16.4 32.6
Frozen Seg Chen et al. (2024b) yes 3.41 / 1.07 4.48 25.9 16.4 34.4
SAM-CP yes 3.41 / 0.21 3.62 27.2 17.0 31.8
SAM2-CP yes 1.61 / 0.18 1.79 27.9 18.0 32.6
Table 11: Inference time comparison. ‘SAM / non-SAM’ means the SAM and non-SAM modules.
Adding extra proposals. We add the proposals extracted by the pre-trained Mask DINO (MD) model into the pool of candidate patches. Details are provided in Appendix B.1. Table 10 shows improved segmentation results, inspiring us that SAM does not generate ideal patches.
@@
Time comparison. We chose two SOTA methods to compare their time performance within our open-vocabulary setting in Table 11. All experiments were carried out using an RTX 4090. Compared with SAM-based FrozenSeg, our method has less time in non-SAM modules. When replacing SAM with SAM2, the speed becomes faster while the performance is higher.
Part segmentation. We conduct part segmentation on the Pascal Part datasets Chen et al. (2014), featuring 4465 images with 93 part categories. Table 10 presents the performance, and Figure 8 offers visualizations. The results’ first and second lines stem from models trained solely on part categories and on both part and general ones, respectively. The model demonstrates stable and reliable part - segmentation performance across different segmentation types. SAM relies on low-level cues like color instead of semantics for predictions, as seen in its tendency to separate arm cuffs from skin. Our SAM-based approach inherits these drawbacks, but future tuning on this data may improve results. Additionally, we crop the original image using the instance segmentation map or bounding box, enlarge the sub-image for extra SAM segmentation to obtain more parts, shown in Figure 7, and iteratively use prompt I and prompt II for finer semantic and instance segmentation.
Figure 5: Direct classification for parts and ’parts of the whole’.
Figure 6: Dynamic prompts for looking for small objects.
Figure 7: Interactive SAM calling yields finer results.
Figure 8: The visualization of part and general instance segmentation. The result is obtained with one model and text labels of different granularities. On the left are the GTs, and on the right are the results.
4.6 Discussion
Advantages. SAM-CP inherits a clear advantage from SAM in the generalized ability across different visual domains. SAM-CP achieves this by decomposing low-level pixel grouping (offered by SAM) from high-level semantic recognition (offered by the composable prompts). In the meantime, the imperfection of SAM limits the segmentation accuracy of SAM-CP, especially in the closed-set vision benchmarks. We showcase this point in Table 10 where the patches generated by SAM are compared to those generated by Mask DINO Li et al. (2023b), a state-of-the-art segmentation model. We find that SAM suffers a higher missing rate (e.g., one cannot find a proposal with an IoU larger than $0.5$ for $16.7\%$ instances; the rate is not significantly smaller even if we refer to the ground-truth masks to merge some proposals) than Mask DINO (where the rate is only $10.1\%$).
Classification mechanism. SAM-CP does not directly classify each patch but uses a query to conduct cross-attention to all high-affinity patches and then fuses their features as the query feature for classification. This enables the model to use richer and more complete information for semantic classification. Figure 7 shows an example. Three patches are present, corresponding to a dog’s torso, tail, and leg. Both ‘cat’ and ‘dog’ labels are in the semantic space, but it is difficult for each of the three patches alone to classify itself as ‘dog’ or ‘cat’. The model, instead of assigning a high affinity between a class label (‘dog’ or ‘cat’) and any patch, assigns a high affinity between these three patches (note that, no matter what the class is, ‘dog’ or ‘cat’, they belong to the same instance). After they have been merged, the visual features are often sufficient for classification (‘dog’ vs. ‘cat’). In the example, the ‘dog’s leg’ patch gets an ambiguous classification score (dog: 0.4844, cat: 0.5156) initially, but when it is merged with the other patches (as a ‘dog’), the score becomes distinguishable.
Limitations. In other words, SAM failed to find some objects or incorrectly merged two or more objects into one patch, and SAM-CP cannot save such loss. This contributes to the deficit in the segmentation accuracy (e.g., AP or mIoU) compared to Mask DINO. We display some typical examples in Appendix A.3. This issue can be alleviated by dynamically adding denser point prompts to the regions (e.g. small objects). We show a case in Figure 7, where small targets were found with this simple mechanism. Additionally, the inference speed of SAM-CP is bound by that of SAM; once a more efficient vision foundation model is available, our framework can be seamlessly transplanted.
@@ 5 Conclusions
Acknowledge
This work was supported in part by the Key Deployment Program of the Chinese Academy of Sciences, China under Grant KGFZD145-23-18 and the Strategic Priority Research Program of Chinese Academy of Sciences under Grant E1XA310103. In addition, we would like to thank Zhaoyang Wei for his technical support during the article - experiment process and the rebuttal phase.
@@ A.1 Affinity Computation
Dynamic cross-attention. Rather than performing global computation, we use dynamic cross-attention (DCA) to guide the cross-attention operator in focusing on the local features. This is motivated by deformable attention Zhu et al. (2021). The cross-attention operator follows the same structure of multi-head attention Carion et al. (2020), and the matrix $\mathbf{A}$ serves as the dynamic attention mask. We binarize $\mathbf{A}$ by setting the value of each entry to $1$ if the value is smaller than the threshold (e.g., $0.5$) which means that the area is masked and ignored in computing the cross-attention. As described in Section 3.2.2, the query vectors (as Q) and the patch features (as both K and V) are fed into a multi-head cross-attention with the dynamic attention mask.
@@ A.2 Supervision
Ground-truth class labels. Let $c_{m}^{\star}$ be the ground-truth class label for the $m$-th query. For a Type-I prompt, the $m$-th query corresponds to the $c_{m}^{\star}$-th category label. Hence, the objective of the $m$-th query is set to be $c_{m}^{\star}$ if the image has the $c_{m}^{\star}$-th category label, and ‘negative’ otherwise. For a Type-II prompt, $c_{m}^{\star}$ comes from the label assignment of the Hungarian matching algorithm. The matrix comes from the classification cost, mask focal loss cost, and dice cost. Hence, the objective of the $m$-th query is set to be $c_{m}^{\star}$ if it is a positive (matched) query, and ‘negative’ otherwise.
@@
The bottom half of Figure 9 shows: (1) The left part is the category assignment processing. We use the Hungarian Matching to achieve one-to-one matching between queries and GTs. The matrix comes from the classification cost, mask focal loss cost, and dice cost. (2) The middle part is the t-SNE visualization between GTs (GT1, GT2, …) and patches. After the category assign procedure, some queries are assigned to the GTs while others are not matched. (3) the right part is the t-SNE visualization which shows the relationship between patches and queries (Q1, Q2, …).
Figure 9: The illustration of how to get the ground-truth affinity matrix $\mathbf{B}$. The left is GTs&\&&Q, the middle is GTs&\&&P and the right is Q&\&&P. Line 2 is the t-SNE visualization of category assignment.
@@
A.3 Examples for Advantages and Limitations
@@
Data preparation. We use the same data augmentation strategy in DETR-based detector Carion et al. (2020); Zhu et al. (2021); Zhang et al. (2023a). We randomly flip the image, resize the image ranging from $1333\times 480$ to $1333\times 800$, randomly crop the image, and randomly resize the image again. In the inference stage, we resize the image to $1333\times 800$ without any other augmentation. For SAM proposal generation, we use the everything mode in SAM with $48$ points per side, a $0.8$ IoU threshold, a $0.9$ stability score threshold, and a $0.7$ NMS threshold.
@@
Add patches from Mask DINO. We add the patches from Mask DINO Li et al. (2023b) to show that the upper bound of SAM patches limits our performance. If the quality of patches generated by the foundation model gets better in the future, the segmentation accuracy of SAM-CP will further increase. We saved the segmentation results of Mask DINO and deleted the label result and confidence score. Then, the segmentation results are viewed as the extra proposal and added to the SAM patches.
@@
Data preparation. We use the same data augmentation strategy in Mask2former Cheng et al. (2022) in ADE20K dataset. The image is randomly flipped and resized in the range of $[320,1280]$. Then, the image is randomly cropped by $2560\times 640$. In the inference stage, the image is resized to $2560\times 640$. The hyper-parameters for SAM patch generation are the same as those for the COCO dataset.
@@
Setting for the model. The training hyper-parameters are the same as those in close-domain experiments. Only the trainable backbone is replaced by a frozen CLIP image encoder. In the inference stage, the parameter $\kappa$ described in Section 4.1 is set as 0.4. In addition, we follow FCCLIP Yu et al. (2023a) to use the prompt engineering when inference. Each label can derive multiple entries, and we chose the highest score as the classification score for this label.
@@
We provide our codes in code.zip file in the Supplementary Material. After the acceptance of our paper, we will make our codes publically available.
@@
Appendix C More Experiments
C.1 More Ablation Studies
C.2 More Comparison
Comparison with SAM-based method. Then, we compared our method with OVSAM [1] and prompt segment anything [3] on the COCO dataset to evaluate the instance segmentation ability on Table 13. All other methods need a pre-trained detector on COCO, and the detection results are used as the prompt for SAM. But our localization only relies on the original SAM, without tuning on COCO. The results show our performance is the best:
$\tau$ PQ APdet AP mIoU
0.9 43.2 42.8 38.2 49.9
0.8 45.0 45.4 40.7 52.6
0.7 43.3 43.3 39.9 50.3
0.5 42.0 34.3 37.2 48.9
0.8 (w/o $\mathrm{IoP}_{\mathrm{box}}$) 42.1 31.7 35.4 48.4
0.8 (w/o $\mathrm{IoP}_{\mathrm{mask}}$) 39.7 42.3 32.5 44.1
0.8 (w/low quality match) 47.0 45.8 41.4 54.2
Method Backbone Detector mAP AP50
OVSAM [a] R50 FRCNN 35.8 55.6
OVSAM [a] Swin - B Detic 36.7 57.2
Prompt segment anything [b] Res50 H - Deformable - DETR+tricks 41.5 -
SAM - CP(ours) R50 41.7 59.4
Table 13: Instance segmentation comparison with SAM-based method.
The comparison between our method and SEEM (Table 14) on the COCO dataset is listed as follows. With the same amount of backbone, we achieved higher performance. We also added a comparison with Semantic SAM (Table 15). Even though it has point prompts in inference, our method outperforms them by a lot, which fully demonstrates that our method is superior to other SAM-based methods.
Method Backbone PQ mAP mIoU
SEEM [c] ViT - L 52.0 43.5 60.2
SAM - CP(ours) Swin - L 52.7 45.2 61.8
Table 14: Comparison of panoptic segmentation metrics with SEEM.
Method Backbone Prompt(testing) mIoU
Semantic - SAM [d] Swin - L point 57.0
Semantic - SAM [d] (reproduce) Swin - L point 55.1
SAM - CP(ours) Swin - L None 61.8
Table 15: Comparison of semantic segmentation with different SAM-based methods.
By iteratively calling our composable prompts, SAM-CP achieves the instance and semantic segmentation based on the vision foundation model SAM without mask tuning. For SAM2-CP in the first table of this question, we substitute SAM with SAM2 [5] in SAM-CP. SAM2 is faster than SAM. We utilize the CP trained on SAM-CP and transfer the weights to SAM2-CP without re-train, with the performance being 27.9 PQ. This indicates that our method is modular and can be extended to different patch generation methods (whether stronger or faster). Although the current SAM-based method is more time-consuming than the non-SAM-based method, with the improvement of the segmentation accuracy and effectiveness of the SAM-based foundational model, the new bottom-up perception paradigm designed based on its high generalization ability is valuable, which is also one of our main motivations.
[a] Open-Vocabulary SAM. ECCV 2024
[b] Prompt segment anything,(https://github.com/RockeyCoss/Prompt-Segment-Anything)
[c] Segment Everything Everywhere All at Once. NeurIPS 2023
[d] Semantic-SAM: Segment and Recognize Anything at Any Granularity, arXiv:2307.04767
Traing time comparison. In addition, SAM-CP only needs 12 epochs of training, while 50 epochs are required for FCCLIP and FrozenSeg.
Appendix D More Visualization Results
D.1 Visualization of cross-domain experiments
For out-of-domain generalization, we conducted experiments on camouflage object segmentation and provided some visualizations in Figure 12.
D.2 More t-SNE Visualization
Figure 13 shows more t-SNE visualizations of our approach. The features of patches belong to different semantic/instance groups to different clusters in SAM-CP. Not that, in lines 6–10, the different instances with the same label (‘train’, ‘elephant’, ‘person’, ‘horse’, and ‘umbrella’) cluster together to the same semantic area but separate to different instances in the t-SNE feature space.
D.3 More Segmentation Examples